# Optimal Best-Arm Identification under Fixed Confidence with Multiple Optima

### v1

@@ VIII Conclusion
In this work, we revisited the problem of best-arm identification in stochastic multi-armed bandits under the fixed-confidence setting, with a particular focus on instances where multiple optimal arms may exist. While the Track-and-Stop algorithm has been widely regarded for its strong empirical performance, its theoretical guarantees in this more general setting have remained an open question.
Our primary contribution is a rigorous proof that Track-and-Stop achieves instance-optimal sample complexity even when the set of optimal arms is not a singleton. By carefully analyzing the generalized likelihood ratio statistics and aligning the stopping condition with information-theoretic lower bounds, we demonstrate that the algorithm retains its optimality across all problem instances—without requiring any modifications to its original formulation.
This result strengthens the theoretical foundation of likelihood ratio-based strategies for best-arm identification and affirms the robustness of Track-and-Stop beyond the classical unique-optimum setting. Looking ahead, an important direction for future work is to extend these insights to more structured bandit models, such as combinatorial or contextual bandits, and to develop adaptive algorithms that offer both optimal guarantees and improved empirical performance.
@@ Appendix A Proof of Lemma 2
$d(\mu,\lambda)=D(P_{\mu}\|P_{\lambda})=A(\theta(\lambda))-A(\theta(\mu))-(\theta(\lambda)-\theta(\mu))\mu.$ (95)
@@ Appendix B Proof of Lemma 7
Now, since $A\geq\log e$, we have $-e^{-A}\in[-\frac{1}{e},0)$. This means that
@@
Now, by (Chatzigeorgiou2013BoundsOT, ), it holds that
$-1-\sqrt{2u}-u<W_{-1}(-e^{-u-1})<-1-\sqrt{2u}-\frac{2}{3}u,\qquad\forall u>0.$ (102)
It follows that
@@ Acknowledgements
The authors sincerely thank Jonathan Scarlett for his valuable suggestions that helped improve the manuscript.

### v2

@@ VIII Conclusion
In this work, we revisited the problem of best-arm identification in stochastic multi-armed bandits under the fixed-confidence setting, focusing on the general case where multiple optimal arms may exist. Unlike prior studies, which primarily addressed the scenario with an unknown number of optimal arms, we investigated the setting where this number is known in advance.
Our main contribution lies in deriving a new information-theoretic lower bound that characterizes the minimal achievable sample complexity when the number of optimal arms is known. Building on this insight, we established that a modified Track-and-Stop algorithm—equipped with an appropriate stopping rule based on generalized log-likelihood ratio statistics—achieves this bound asymptotically. This result provides the first instance-optimal guarantee for Track-and-Stop in multi-optimal settings and demonstrates that knowledge of the number of optimal arms can fundamentally reduce sample complexity.
Overall, our findings extend the theoretical foundations of best-arm identification and reinforce the role of likelihood ratio-based methods as principled and optimal strategies for sequential decision-making under uncertainty. Future research directions include generalizing these results to structured bandit models, such as combinatorial or contextual settings, and developing adaptive algorithms that leverage structural information while preserving optimality guarantees.
@@ Appendix A Proof of Lemma 2
$d(\mu,\lambda)$
$\quad=D(P_{\mu}\|P_{\lambda})$ (128)
$\quad=A(\theta(\lambda))-A(\theta(\mu))-(\theta(\lambda)-\theta(\mu))\mu.$ (129)
@@ Appendix B Proof of Lemma 7
Now, since $A\geq\log e$, we have $e^{-A}\in[-\frac{1}{e},0)$. This means that
@@
Now, by [14], it holds that
$-1-\sqrt{2u}-u<W_{-1}(-e^{-u-1})<-1-\sqrt{2u}-\frac{2}{3}u$ (136)
for all $u>0$. It follows that
@@ Acknowledgements
The author would like to thank Ho Chi Minh City University of Technology (HCMUT), Vietnam National University, Ho Chi Minh City (VNU-HCM), for supporting this study. The author also sincerely thanks Jonathan Scarlett for his insightful suggestions, which significantly improved the manuscript. Appreciation is further extended to the Associate Editor and the reviewers for their constructive comments and valuable feedback, which helped enhance the overall quality of the paper.